DIFFUSION OF HYDROGEN FLUORIDE IN SOLID PARAHYDROGEN.

\underline{H. OOE}, Y. MIYAMOTO, J. TANG, K. KAWAGUCHI, Graduate School of Natural Science and Technology, Okayama University, 3-1-1 Tsushima-naka Okayama 700-8530,Japan; S. KUMA, K. NAKAJIMA, N. SASAO, T. TANIGUCHI, Research Core for Extreme Quantum World, Okayama University, 3-1-1 Tsushima-naka Okayama 700-8530,Japan; I. NAKANO and M. YOSHIMURA, Faculty of Science, Okayama University, 3-1-1 Tsushima-naka Okayama 700-8530,Japan.

In general, atoms and molecules diffuse thermally in solid with rate which has Arrhenius-type temperature dependence. On the other hand, it is known that diffusion rate at low temperature sometimes shows non-Arrhenius behavior, which is called gquantum diffusionh^{a}. We have studied hydrogen fluoride (HF) in solid parahydrogen (p-H_{2}) by FTIR absorption spectroscopy and found that HF diffuses in solid p-H_{2} even at 3.6 K^{b}. In this study, dependence of the rate on temperature, HF concentration and IR irradiation was investigated. Assuming Arrhenius-type dependence, activation energy of the diffusion is less than a few Kelvin, which suggests that the diffusion has the quantum nature. Recent experimental results and a possible mechanism of the diffusion will be discussed.

aY. Kagan and A. J. Leggett gQuantum Tunneling In Condensed Mediah, North Holland, 1992.

bY. Miyamoto et al. J. Phys. Chem. A \underline{115}, 14254 (2011).